For which values of A , B , C and k is the function yp(x) = x 2 (A x2 + B x + C) a particular solution of the equation 16 y 000 + 9 y 00 = 1458 x k−1 (2x − 1) ? (a) A = 27 , B = 165 , C = −880 , k = 2 (b) A = 27 , B = −219 , C = 1168 , k = 2 (c) A = 27 , B = −219 , C = 1168 , k = 4 (d) A = 54 , B = −411 , C = 2192 , k = 3 (e) A = 54 , B = −411 , C = 2192 , k = 4 


1
Expert's answer
2022-01-10T15:16:42-0500
"16y'''+9y''=1458 x^{ k\u22121} (2x \u2212 1)"

Corresponding homogeneous differential equation


"16y'''+9y''=0"

Characteristic (auxiliary) equation


"16r^3+9r^2=0"

"r^2(16r+9)=0"

"r_1=r_2=0, r_3=-\dfrac{9}{16}"

The general solution of the homogeneous differential equation is


"y=c_1+c_2x+c_3e^{-(9\/16)x}"


The particular solution of the non homogeneous differential equation is


"y_p=Ax^4+Bx^3+Cx^2"

"y_p'=4Ax^3+3Bx^2+2Cx"

"y_p''=12Ax^2+6Bx+2C"

"y_p'''=24Ax+6B"

Substitute


"16(24Ax+6B)+9(12Ax^2+6Bx+2C)"

"=1458 x^{ k\u22121} (2x \u2212 1)"

"x^2=x^{k-1+1}"

"108A=2916"

"384A+54B=-1458"

"96B+18C=0"

"k=2, A=27, B=-219, C=1168"

(b)


"A=27, B=-219, C=1168, k=2"